OpenStudy (anonymous):

The segment AB is divided into 4 equal parts.C is nearer to A and E is nearer to B.Find the coordinates of A and B ,if the co-ordinates of C, D and E are [5/2,-1/2],(3,0) and [7/2,1/2] respectively.

OpenStudy (zehanz):

C, D and E being the points that divide AB into the 4 equal parts, I guess. Let's say A=(a1,a2) and B=(b1,b2). If you want to calculate the midpoints of two given points, you have to calculate the mean of the coordinates: e.g. D is the midpoint of CE, so D=((2.5+3.5)/2, (-.5+.5)/2)=(3,0). Right! The same calculation can be used for A and D: C is midpoint, so the coordinates of C should be: c1=(a1+3)/2=5/2 c2=(a2+0)/2=0 This gives you A. Do the same with B!

OpenStudy (zehanz):

@singhmmm: Did you get any further with this one?

OpenStudy (anonymous):

a1=2 and a2=-1 thus the cordiantes of A =2,-1 is it right @ZeHanz

OpenStudy (anonymous):

simiallry, thus the cordiantes of B=4,-1

OpenStudy (anonymous):

oops, coordiantes of B are 4,1

OpenStudy (zehanz):

a1=2: OK a2=-1... I have: (a2+0)/2=0, so a2+0=0, so a2=0 b1=4: OK b2=-1... I have: (b2+0)/2=1/2, so b2+0=1, so b2=1

OpenStudy (zehanz):

Easy to make mistakes with this calculation ;)

OpenStudy (anonymous):

but plse see a2, its (a2+0)/2=-1/2

OpenStudy (zehanz):

You're right, easy for me too ;)

OpenStudy (anonymous):

but, its okay.. no pblm thank u so much

OpenStudy (zehanz):

yw!